# Forecasting Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis

### v1

@@ Abstract
The forecasting of multivariate urban data presents a complex challenge due to the intricate dependencies between various urban metrics such as weather, air pollution, carbon intensity, and energy demand. This paper introduces a novel multivariate time-series forecasting model that utilizes advanced Graph Neural Networks (GNNs) to capture spatial dependencies among different time-series variables. The proposed model incorporates a decomposition-based preprocessing step, isolating trend, seasonal, and residual components to enhance the accuracy and interpretability of forecasts. By leveraging the dynamic capabilities of GNNs, the model effectively captures interdependencies and improves the forecasting performance. Extensive experiments on real-world datasets, including electricity usage, weather metrics, carbon intensity, and air pollution data, demonstrate the effectiveness of the proposed approach across various forecasting scenarios. The results highlight the potential of the model to optimize smart infrastructure systems, contributing to energy-efficient urban development and enhanced public well-being.
@@ 1 Introduction
Analyzing patterns and predicting trends in urban data, such as electricity demand Nti et al. (2020), air quality Huang et al. (2021); Du et al. (2024), weather and climate Mouatadid et al. (2024), and carbon intensity Maji et al. (2022), is critical for a wide range of applications, from infrastructure planning and operation to disaster prevention, to addressing social inequalities and improving quality of life. Urban data is normally captured by a number of multimodal sensors installed in a large geographical area. Due to the diverse nature of environments in which the sensors are installed, variations in sensor types and specifications, and complex spatial-temporal dependencies within the data, designing accurate prediction models that can be transferred to various domains is not trivial.
@@
Although multivariate time-series forecasting has been extensively studied in the literature, achieving accurate long-term predictions–specifically for horizons ranging from 2 to 6 weeks–remains a significant challenge compared to short- and medium-term forecasting. This difficulty arises from increased uncertainty, changing dynamics, and accumulation of errors. Furthermore, when data is sourced from multiple sensors with diverse modalities, there is currently no universally effective solution that consistently outperforms other models in different domains.
Physics-based dynamical models are either unavailable or inaccurate for long-term predictions. Traditional statistical methods, recurrent neural networks, and Transformer-based models often fail to capture the complexities of multivariate time-series data, leading to large errors in long-term forecasts. In recent years, foundational models have shown promise in time-series forecasting due to their ability to model complex temporal dependencies and handle multivariate data. However, their effectiveness depends on data availability and they may not generalize well to different time-series data, requiring extensive fine-tuning in some cases. Graph Neural Networks (GNNs) offer a promising solution by effectively modeling temporal and spatial dependencies and improving forecast accuracy Wu et al. (2020). This paper explores the application of GNNs to address the challenges of long-term multivariate time-series forecasting in domains such as energy and transportation. Using these advanced models, we aim to improve the accuracy and reliability of the forecast, where reliability refers to the consistency and robustness of predictions across varying conditions. This, in turn, facilitates the optimization of smart infrastructure systems, which are crucial for developing energy-efficient cities and improving public well-being.
Among GNN models, the Graph Attention Network (GAT) Veličković et al. (2017) introduces an attention mechanism to compute edge weights between neighboring nodes based on node features. This improves neighbor information aggregation and enhances GNN performance in various tasks. However, once trained, GAT’s attention mechanism remains fixed, limiting its ability to adapt to evolving graph structures or changing relationships. Building on this foundation, an improved variant called GATv2 Brody et al. (2021) was developed to address the limitations of the original GAT. GATv2 introduces a more expressive attention mechanism that conditions attention scores on both the query and key nodes, enabling it to better capture complex node interactions. Although this enhancement improves flexibility during training, the learned attention weights of the model remain unchanged after training, similar to GAT. This makes GATv2 particularly effective for tasks requiring richer relational modeling, such as multivariate time-series forecasting.
We propose a novel multivariate time-series forecasting model that takes advantage of GATv2’s power to learn spatial dependencies among the variates and Temporal Convolutional Network (TCN) Bai et al. (2018) to learn temporal dependencies within each variate. Our approach involves constructing a graph where nodes represent different variates and edges represent relationships inferred from historical data in an offline mode. Using GATv2, our model dynamically adjusts the attention given to each node’s neighbors, effectively capturing interdependencies and improving forecast accuracy. We further enhance our model by incorporating a pre-processing step that decomposes the time-series data into its fundamental components: trend, seasonal, and residual. This decomposition allows our model to focus on the inherent characteristics of the data, facilitating more accurate forecasts. Extensive evaluation carried out on different urban datasets demonstrates the superior performance of our approach compared to state-of-the-art time-series forecasting models, highlighting its potential for practical applications in various domains.
@@
  • We propose a novel multivariate time-series forecasting model that uses GATv2 and TCN to learn the salient spatial and temporal features that enhance the accuracy of long-term forecasts.
  • We implement a decomposition-based preprocessing step to isolate the trend, seasonal, and residual components of the time series, improving forecast accuracy.
  • We validate our model’s effectiveness through extensive evaluation on diverse real-world datasets, showcasing significant improvements over the state-of-the-art in time-series forecasting across these datasets.
It is worth noting that, in some applications, forecasts related to multiple types of urban data are often integrated into optimization or decision-making problems to minimize costs, energy consumption, and carbon emissions. For example, forecasts of electricity consumption, weather, and carbon intensity must be incorporated simultaneously to determine how to distribute computing jobs across data centers located in different geographical areas, thereby minimizing costs and carbon emissions Radovanović et al. (2023). Thus, even a slight improvement in the prediction accuracy in each dataset could translate to considerable savings in applications that rely on multiple datasets, highlighting the importance of this contribution.
@@ 2 Related Work
Classical statistical methods have been the cornerstone of time-series forecasting for decades. Models such as the Autoregressive Integrated Moving Average (ARIMA) Box and Jenkins (1968); Box and Pierce (1970) and its variants have historically been used due to their simplicity and interpretability. ARIMA models work by modeling the autocorrelation in the data, making them suitable for univariate time-series forecasting. Another well-known method is Exponential Smoothing (ETS) Holt (2004); Winters (1960), which assigns exponentially decreasing weights to past observations. Despite their widespread use, these classical methods have notable limitations. They typically assume stationarity and linearity in the data, which can make them unsuitable for more complex, nonlinear, and large-scale multivariate time-series datasets. Furthermore, as the number of variables increases, these models often struggle with scalability and may overfit data. This underscores the need for more advanced techniques that can handle the intricacies of urban time-series data.
@@
Machine learning has brought new techniques to time-series forecasting. Traditional models like Support Vector Machines (SVM) Boser et al. (1992), Decision Trees, and Random Forests Breiman (2001) have been adapted to predict future values based on historical data. These models are capable of capturing nonlinear relationships better than classical statistical methods. Ensemble methods such as Gradient Boosting Machines (GBM) Friedman (2001) and XGBoost Chen and Guestrin (2016) have been particularly successful in combining the predictions of multiple models to improve accuracy. However, these models often require significant feature engineering and may not fully capture the temporal dependencies inherent in time-series data.
@@
Deep learning has revolutionized time-series forecasting by making it possible to learn nonlinear and complex temporal dependencies directly from the data. Early deep learning models for time-series forecasting, such as Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) networks and Gated Recurrent Units (GRUs) Chung et al. (2014), focused on capturing long-term dependencies through recurrent neural network (RNN) architectures.
@@
More recent advances include models such as N-BEATS Oreshkin et al. (2019), which employs backward and forward residual links and a deep stack of fully connected layers to achieve state-of-the-art performance on various benchmark datasets. Transformer-based models, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and Fedformer Zhou et al. (2022), have also shown significant promise by leveraging self-attention mechanisms to capture long-range dependencies more effectively than RNN-based approaches. However, some recent studies show that transformer-based models may not always be the best fit for time-series forecasting Zeng et al. (2023). Transformer-based models may struggle with effectively capturing the inherent temporal patterns in time-series data due to their reliance on self-attention mechanisms. DLinear and NLinear models Zeng et al. (2023) take a simpler and more efficient approach that mitigates these issues by focusing on linear models that are more adept at handling the sequential nature of time-series data.
@@
One recent related work is CycleNet Lin et al. (2024), which introduces Residual Cycle Forecasting (RCF) to model periodic patterns using learnable recurrent cycles and predict residual components. However, CycleNet assumes stable periodicities, making it less effective for dynamic or irregular cycle datasets. Additionally, it models each time-series channel independently, failing to capture interdependencies between variables, which can limit its performance on multivariate time-series data.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for time-series forecasting, particularly for multivariate time-series data. GNNs are designed to capture spatial dependencies by representing the data as graphs, where nodes represent variables, and edges represent dependencies between these variables. One notable approach is the MTGNN Wu et al. (2020) framework, which integrates temporal and graph convolution modules to capture both temporal and spatial dependencies, using a graph learning layer to adaptively construct an adjacency matrix that reflects hidden relationships among time-series variables. Another example is the Multivariate Time-Series Graph Ordinary Differential Equation (MTGODE) Jin et al. (2022) framework, which uses graph neural networks to model complex interactions among time-series and employs neural ODEs to continuously evolve the data over the graph structure, enhancing forecasting accuracy and robustness.
Defining optimal graph structures and integrating graph learning with time-series forecasting in an end-to-end manner remains a significant challenge. While recent advances have introduced adaptive graph learning Zhang et al. (2020) and automated spatio-temporal fusion techniques Li and Zhu (2021), these methods come with their own limitations. Specifically, adaptive graph learning can be computationally expensive, and spatio-temporal fusion mechanisms may struggle to effectively capture long-range dependencies, leading to potential inaccuracies in forecasting Kumar et al. (2024).
@@
Recent advancements in time-series forecasting feature foundation models such as TimesFM Das et al. (2023) and MOIRAI Woo et al. (2024), which generalize across diverse datasets without additional training. TimesFM, using a decoder-only architecture with input patching, achieves near state-of-the-art zero-shot performance. MOIRAI, with a masked encoder architecture, addresses challenges like cross-frequency learning and any-variate forecasting, leveraging the Large-scale Open Time-Series Archive (LOTSA) Woo et al. (2024) for training. Both models demonstrate the potential of foundation models in handling complex and varied time-series forecasting tasks. However, these models can be computationally intensive, require significant resources for training and deployment, and often lack interpretability.
Overall, while each category of time-series forecasting methods has distinct limitations, the integration of graph neural networks with deep learning techniques represents a promising direction to advance the state-of-the-art in multivariate time-series forecasting.
3 Problem Formulation
Let $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ represent a time-series of $T$ observations, each being $d$-dimensional. Time-series forecasting concerns predicting future values based on past observations. Given a look-back window of length $l$ ending at time $t$, denoted as $X_{t,l}=(x_{t-l+1},\ldots,x_{t-1},x_{t})$, the goal is to forecast the next $h$ steps of the time-series. This is achieved by learning the function $f_{\theta}(X_{t,l})=(\hat{x}_{t+1},\ldots,\hat{x}_{t+h})$, where $\theta$ represents the parameters of the forecasting model. We refer to each pair of look-back and forecast windows as a sample.
@@ 4 Methodology
We model the input and output time-series as a graph. Each sensor channel $i$ is represented as a node $v_{i}$ that might have relationships with other channels, represented as directed and weighted edges $e_{ij}$ connecting these nodes. These connections are assumed to be static, i.e., the graph structure $G=(V,E)$ does not change over time. The node features are initially the historical values of the corresponding channels $x_{i}=\{x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{T}\}$, and ultimately the forecast values $\hat{y}_{i}=\{\hat{y}_{i}^{T+1},\hat{y}_{i}^{T+2},\ldots,\hat{y}_{i}^{T+H}\}$.
The proposed methodology has two main parts:
@@
    Graph Structure Learning: This involves learning the adjacency matrix $A$ capturing the relationships between nodes. The matrix $A\in\mathbb{R}^{D\times D}$ where $D$ is the number of nodes, and each element $a_{ij}$ represents the weight of the edge from node $v_{i}$ to node $v_{j}$.
@@
    Spatio-Temporal Forecasting: This is the core of our model, responsible for learning and prediction. It comprises three subcomponents:
    1. (a)
      Decomposition: Each input sample is decomposed into its fundamental components to better capture underlying patterns.
    2. (b)
      Spatio-Temporal Feature Extraction: This step extracts features from the graph using both spatial and temporal information. For each node, $v_{i}$, the feature vector $h_{i}$ is derived by combining its historical values $x_{i}$ and the influence of its neighboring nodes as defined by $A$.
    3. (c)
      Forecasting: Finally, the model uses the extracted features to predict future values. The forecast values $\hat{Y}$ for all nodes are given by $\hat{Y}=f(H,A)$, where $H$ is the matrix of all node features and $f$ is the forecasting function.
Each of these parts will be explained in more detail in the subsequent sections.

### v2

# Long-Term Forecasting of Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis

@@ Abstract
Long-term forecasting of multivariate urban data poses a significant challenge due to the complex spatiotemporal dependencies inherent in such datasets. This paper presents DST, a novel multivariate time-series forecasting model that integrates graph attention and temporal convolution within a Graph Neural Network (GNN) to effectively capture spatial and temporal dependencies, respectively. To enhance model performance, we apply a decomposition-based preprocessing step that isolates trend, seasonal, and residual components of the time series, enabling the learning of distinct graph structures for different time-series components. Extensive experiments on real-world urban datasets—including electricity demand, weather metrics, carbon intensity, and air pollution—demonstrate the effectiveness of DST across a range of forecast horizons, from several days to one month. Specifically, our approach achieves an average improvement of 2.89% to 9.10% in long-term forecasting accuracy over state-of-the-art time-series forecasting models.
@@ 1 Introduction
Analyzing patterns and predicting trends in urban data, such as electricity demand Nti et al. (2020), traffic flow and air quality Huang et al. (2021); Du et al. (2024), weather and climate variables Mouatadid et al. (2024), and carbon intensity Maji et al. (2022), is critical for a wide range of applications, including infrastructure planning, disaster preparedness, and addressing social inequalities. Urban data is typically multimodal and collected across large geographical areas. Due to variations in sensor types and deployment environments, and complex spatiotemporal dependencies within the data, building accurate and transferrable forecasting models is a non-trivial task.
@@
While multivariate time-series forecasting has been widely studied, achieving accurate long-term predictions, specifically over horizons of two to six weeks, remains a significant challenge Zhou et al. (2021). This difficulty arises from the accumulation of errors and shifting dynamics. Furthermore, when data originates from multiple heterogeneous sensors, there is currently no universally effective model that consistently performs well across domains. Physics-based dynamical models, where available, are inaccurate for long-term forecasting Mouatadid et al. (2024). Traditional statistical methods, recurrent neural networks, and transformer-based models often fail to capture the complex dependencies inherent in multivariate time-series data, leading to performance degredation on long-term forecasts.
In recent years, foundational models have shown promise in time-series forecasting due to their ability to model complex temporal dependencies and handle multivariate data. However, their success hinges on access to large-scale pretraining data Woo et al. (2024), and they often require extensive fine-tuning to adapt to new domains, limiting their practical applicability.
Graph Neural Networks (GNNs) offer a promising alternative by naturally capturing both temporal and spatial dependencies, thereby improving forecast accuracy in multivariate settings Wu et al. (2020). However, the underlying graph structure is typically unknown and must be inferred from the data. In this paper, we explore the use of GNNs to address the challenges of long-term forecasting of multivariate urban time-series data, with an emphasis on improving forecast accuracy. A key insight of our approach is that the latent graph structures corresponding to different time-series components (e.g., trend, seasonality) are often distinct. This observation motivates the application of time-series decomposition prior to graph structure learning.
We propose a novel model that combines the Graph Attention Network (GAT) Veličković et al. (2017) for learning spatial dependencies among variates with a Temporal Convolutional Network (TCN) Bai et al. (2018) to capture temporal dependencies within each variate. We construct a graph where nodes represent individual variates and edges encode dependencies inferred from historical data. Leveraging GATv2 Brody et al. (2021), the model dynamically adjusts the influence of each node on neighboring nodes. To further enhance accuracy, we introduce a preprocessing step that decomposes the time-series into its fundamental components: trend, seasonal, and residual components. This enables the model to focus on the intrinsic structure of the data, using a distinct graph structure for each component.
We evaluate our model on multiple real-world urban datasets, demonstrating significant improvements in long-term forecasting performance over the state-of-the-art. These results underscore the model’s potential for practical deployment in diverse application domains.
@@
  • We propose a novel multivariate time-series forecasting model, called DST, that integrates GATv2 and TCN to learn the salient spatial and temporal features, thereby enhancing the accuracy of long-term forecasting.
  • We implement a decomposition-based preprocessing step to isolate trend, seasonal, and residual components of the time-series, and learn a graph structure for each component independently. Together, they allow the model to tailor its predictions to distinct temporal dynamics, significantly improving its accuracy.
  • We validate our model’s effectiveness through extensive evaluation on four real-world datasets, showcasing on average 2.89% to 9.10% improvement over the state-of-the-art in time-series forecasting across these datasets.
We note that accurate forecasts across different types of urban data are often integral to decision-making tasks aimed at minimizing cost, energy consumption, and carbon emissions. For example, effectively distributing workloads across geographically dispersed datacenters requires simultaneous predictions of electricity demand, weather, and carbon intensity Radovanović et al. (2023). Thus, even a slight improvement in the prediction accuracy in each dataset could translate to considerable savings in applications that rely on multiple such datasets. This highlights the broader impact of our contributions.
@@ 2 Related Work
Statistical methods have been the cornerstone of time-series forecasting for decades. Time-series models such as the Autoregressive Integrated Moving Average (ARIMA) Box and Pierce (1970) and its variants have historically been used due to their simplicity and interpretability.
@@
Another widely used method is Exponential Smoothing (ETS) Holt (2004); Winters (1960), which assigns exponentially decreasing weights to past observations. Despite their widespread use, these classical methods have notable limitations. They typically assume stationarity and linearity in the data, which can make them unsuitable for more complex, nonlinear, and large-scale multivariate time-series datasets. Furthermore, as the number of variables increases, these models often struggle with scalability and may overfit data.
@@
Traditional machine learning models such as Support Vector Machines (SVM), Decision Trees, and Random Forests have been adapted to predict future values based on historical data. These models are capable of capturing nonlinear relationships better than statistical methods. Ensemble methods such as Gradient Boosting Machines (GBM) Friedman (2001) and XGBoost Chen and Guestrin (2016) have been particularly successful in combining the predictions of multiple models to improve accuracy. However, these models often require significant feature engineering and may not fully capture the temporal dependencies inherent in time-series data.
@@
Deep learning has revolutionized time-series forecasting by making it possible to learn nonlinear and complex temporal dependencies directly from the data. Early deep learning models for time-series forecasting
@@
focused on capturing long-term dependencies through recurrent neural network (RNN) architectures. More recent advances include models such as N-BEATS Oreshkin et al. (2019), which employs backward and forward residual links and a deep stack of fully connected layers to achieve state-of-the-art performance on various benchmark datasets. Transformer-based models, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and Fedformer Zhou et al. (2022), have also shown significant promise by leveraging self-attention mechanisms to capture long-range dependencies more effectively than RNN-based approaches. However, some recent studies show that transformer-based models may not always be the best fit for time-series forecasting Zeng et al. (2023). Particularly, they may struggle with effectively capturing the inherent temporal patterns in time-series data due to their reliance on the self-attention mechanism, and have high computational overhead. DLinear and NLinear models Zeng et al. (2023) take a simpler and more efficient approach that mitigates these issues by focusing on linear models that are more adept at handling the sequential nature of time-series data.
To address the high computational overhead of transformer-based models, PatchTST Nie et al. (2022) segments time-series into patches, significantly improving computational efficiency and forecasting accuracy. Similarly, S-Mamba Wang et al. (2025) employs a selective state-space approach to efficiently capture inter-variate correlations and temporal dependencies with nearlinear complexity, achieving state-of-the-art performance across several benchmarks.
@@
Another recent work, CycleNet Lin et al. (2024), introduces Residual Cycle Forecasting (RCF) to model periodic patterns using learnable recurrent cycles and to predict residual components. However, CycleNet assumes stable periodicities, making it less effective for dynamic or irregular cycle datasets, i.e., datasets with regularly repeating patterns such as daily or seasonal patterns. Additionally, it models each time-series channel independently, failing to capture interdependencies between variables, which can limit its performance on multivariate time-series data.
@@
Recent advances in time-series forecasting feature foundation models such as TimesFM Das et al. (2024) and MOIRAI Woo et al. (2024), which generalize across diverse datasets without additional training. TimesFM, using a decoder-only architecture with input patching, achieves near state-of-the-art zero-shot performance. MOIRAI, with a masked encoder architecture, addresses challenges like cross-frequency learning and any-variate forecasting, leveraging the Large-scale Open Time-Series Archive (LOTSA) Woo et al. (2024) for training. Both models demonstrate the potential of foundation models in handling complex and varied time-series forecasting tasks. However, these models can be computationally intensive, require significant resources for training and deployment, and often lack interpretability.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for multivariate time-series forecasting by modeling variables as nodes and their interdependencies as edges in a graph. The MTGNN Wu et al. (2020) framework introduces a joint architecture combining temporal and graph convolution modules, with a graph learning layer that adaptively constructs an adjacency matrix to uncover hidden variable relationships. In contrast, MTGODE Jin et al. (2022) employs neural ordinary differential equations over dynamically learned graphs to model continuous spatial-temporal dynamics, offering improved accuracy and robustness in forecasting tasks.
Despite the large amount of work that utilizes GNNs for time-series forecasting, discovering optimal graph structures and integrating graph learning with time-series forecasting remains a significant challenge. While recent advances have introduced adaptive graph learning Zhang et al. (2020) and automated spatio-temporal fusion techniques Li and Zhu (2021), these methods come with their own limitations. Specifically, adaptive graph learning can be computationally expensive, and spatio-temporal fusion mechanisms may struggle to effectively capture long-range dependencies, leading to potential inaccuracies in forecasting Kumar et al. (2024).
To evaluate the proposed GNN-based long-term forecasting model, we consider the best model from each of the above categories—namely linear models, state-space models, transformer-based models, and foundation models—as our baseline.
@@ 3 Methodology
Let $X=(x^{1},x^{2},\ldots,x^{T})\in\mathbb{R}^{T\times D}$ represent a time-series of $T$ observations, each being $D$-dimensional, i.e. emitted by $D$ sensors. Time-series forecasting concerns predicting future values based on past observations. Given a look-back window of length $l$ ending at time $t$, denoted as $X^{t,l}=(x^{t-l+1},\ldots,x^{t-1},x^{t})$, the goal is to forecast the next $h$ steps of the time-series. This is achieved by learning the function $f_{\theta}(X^{t,l})=Y^{t,h}=(y^{t+1},\ldots,y^{t+h})$, where $\theta$ represents the parameters of the forecasting model. We refer to each pair of look-back and forecast windows as a sample. Since we focus on long-term forecasting, the forecast window is assumed to be long, usually several days to a few months.
We model the input and output time-series as a graph $G=(V,E)$. Each sensor $i$ is represented as a node $v_{i}$ that might have dependencies with other sensors. These dependencies are modeled using directed, weighted edges connecting the respective nodes. These connections are assumed to be static, i.e., the graph structure does not change over time.
The node features are initially the historical sensor readings with the look-back window, and ultimately the forecast values.
Our methdology has three main steps:
@@
    Decomposition: Decomposing the time-series into fundamental components (trend, seasonal, residual) to accurately capture inherent characteristics of each component.
@@
    Graph Structure Learning: Learning the adjacency matrix $A\in\mathbb{R}^{D\times D}$ (with $D$ being the number of nodes) that captures the relationships between nodes for each time-series component. Each element of $A$, denoted as $a_{ij}$, represents the weight of the edge from node $v_{i}$ to node $v_{j}$.
  3. 3.
    Spatio-Temporal Forecasting via DST: Utilizing the learned graph structure for each component, a GNN extracts spatio-temporal features for predicting future values of the respective component. The predicted future values of different components are then summed to produce the final forecast $Y^{t,h}$.
Each of these steps will be explained in more detail in the subsequent sections.
3.1 Decomposition
The first step in our approach involves decomposing the time-series into its fundamental components: trend, residual, and one or more seasonal components. This decomposition is crucial as it isolates the inherent characteristics of the time-series, thus improving the accuracy of forecasting Sohrabbeig et al. (2023).
The decomposition method adopted here is inspired by the MSTL decomposition technique Bandara et al. (2025). It begins with the application of a weighted moving average filter to the time-series, where the weights for this kernel are derived from the tricube function given below:
Intuitively, it assigns higher weights to nearby data points and lower weights to distant ones. The extracted trend is then subtracted from the original time-series to obtain the detrended series.
Next, the period of seasonality (e.g., daily or weekly patterns) must be determined. This can be done using various methods, such as autocorrelation analysis, Fourier transform, data visualization, etc.
The next step is to capture seasonal patterns in the detrended series. A simple aggregation approach with a sliding window of appropriate length (denoted as $k$) and stride (denoted as $m$) is used for this purpose. The seasonal pattern is identified by averaging the segments covered by the sliding window.
Figure 1: Extracting seasonal patterns using a sliding window of size $k$ and a stride of $m$ times the length of the seasonality period, respectively. In this example, $k=3$ and $m=1$.
Figure 1 illustrates this process. In each step of moving the sliding window, $S_{i}$ is calculated using the following equation:
$S_{i}=\frac{1}{k}\sum_{j=i-\lfloor k/2\rfloor}^{i+\lfloor k/2\rfloor}B_{i},$ (2)
where $B_{i}$ is the i-th segment of the detrended time-series. Each segment contains $p$ data points, where $p$ represents the length of the seasonality period inferred in the previous step. Hence, the data points in $B_{i}$ have indices $i\times p$ to $(i+1)\times p-1$. To preserve the input length, padding is added to both ends of the time-series. This padding consists of copies of the first and last blocks.
Finally, the so-obtained seasonal components are subtracted from the detrended series to obtain the residual component.